# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems

### v1

@@ 9. Discussion
Through our systematic literature review and analysis of real-world CSS implementations, we have produced a theoretical framework that can help the HCI research community better reason about, and design, CCS. These systems were created with the stated goal of addressing the misinformation problem (Coleman, 2021), but it is unclear if the platforms implementing them are serious about this aim. For example, many see Meta’s choice to replace their U.S. fact-checking program with a CCS (Kaplan, 2025) as a thinly-veiled effort to curry political favor with a new presidential administration that is critical of fact-checking (Nix et al., 2025). In fact, from a platform’s perspective, CCS have other appealing properties: they cost less than fleets of fact-checkers, they shield platforms from accusations of biased policy enforcement, and they offload the difficult responsibility of distinguishing fact from fiction onto users. It may be expedient for platforms to build CCS in the current moment, but it is unclear if such systems will survive if the political climate shifts. Still, as long as CCS are in a position to exert such influence over the information ecosystem, it is important for the HCI research community to study them. Most importantly, we emphasize the need to better understand how CCS perform compared to other content moderation systems. CCS may perform well in certain cases, but if on the whole they perform worse than the systems they are replacing, for example on the question of coverage, they should be used carefully.
@@
While it has been suggested that they can, CCS cannot effectively replace a fuller spectrum of content moderation tools, and were not designed to do so (Elliott, [n. d.]). Content moderation is a broad term that encompasses a variety of different goals, including preventing crimes, mitigating lawful but awful behavior, and promoting high quality content. The tools used as a part of content moderation are similarly varied and include content take-downs, labeling, and account bans. Systems that provide context are much narrower in the scope of their goal and mechanism: they aim to better inform users by providing them with additional information. And yet, the empirical results summarized in this paper demonstrate that CCS may play a meaningful role in the content moderation ecosystem, and it is worth exploring how they may integrate with other parts of it. For example, the research that we reviewed found that notes often build on the work done by professional fact-checkers (Borenstein et al., 2025). Might this relationship flow the other way, as well? Perhaps data from CCS systems could be used by professional fact-checkers to cover a wider range of issues. Or, the ability of CCS systems to identify notes with broad appeal could help professionals craft their own fact-checks that do the same.
@@
Our design space reveals several ways forward for a HCI research agenda on CCS. Several aspects are underexplored by current research, particularly the Participation aspect. Encouraging positive participation in collaborative computing systems is a classic topic in HCI research (Kraut and Resnick, 2012) and one in which the community could make strong contributions. One dimension of participation which has not yet been explored is how platforms may play a more active role in soliciting notes and ratings from contributors. Past HCI work on how to best make use of the various skills of crowds may be instructive here (Bernstein et al., 2015; Kittur et al., 2013). It is worth noting that the more consequential these systems become, the more they will be targets for negative forms of participation as well, primarily gaming (Daniel et al., 2018; Epstein et al., 2020). Bridging algorithms are designed to be resistant to certain types of gaming (Wojcik et al., 2022), but vulnerabilities may emerge. One other area for study is the Presentation aspect, specifically how different approaches to visually displaying notes may affect user understanding of misleading claims (Gamage et al., 2025; Martel and Rand, 2023). Another promising area is the Inputs aspect, where recent work has begun to explore the ways in which LLMs can supplement human effort and increase the quality and quantity of notes (De et al., 2025; Mohammadi and Yasseri, 2025). Of particular interest is X’s recent decision to allow AI note contributors via a new API (Li et al., 2025). How such a decision affects the quantity and quality of notes on the platform should be a near-term focus of HCI research. Of course, studying how changes to CCS impact information dynamics across a platform will be difficult to study without more innovations in the Transparency aspect. X’s approach to transparency in their Community Notes system has been a boon for for studying CCS. Hopefully other platforms will soon adopt similar approaches to support ongoing studies of CCS effectiveness.
@@ 10. Conclusion
Crowdsourced Context Systems are an increasingly prevalent part of the content moderation strategies of many prominent platforms. To understand these systems, their impacts, and how they may be better designed, we conducted a systematic literature review of existing research studying implementations of CCS and analyzed several major social media platforms’ implementations of such systems. Despite recent trends towards the adoption of these systems, we emphasize that they are only one piece of a larger content moderation suite, and suggest against using them as drop-in replacements for traditional fact-checking approaches. The framework we produced can be used to reason about these systems, and consists of a theoretical model, a design space, and several key normative concerns. Such systems have potential to address the challenge of misinformation, but to realize this potential, future design research from the HCI community must focus on ways to understand and improve these systems’ performance.

### v2

@@ 8. Discussion
Through our systematic literature review and analysis of real-world CCS implementations, we have produced a theoretical framework that can help the HCI research community better reason about, and design, these systems. In this section, we discuss future directions for CCS research in HCI. Specifically, we focus on how CCS fit within the broader field of content moderation, highlight underexplored portions of the design space, and reflect on the normative implications of CCS design choices.
8.1. Can CCS Replace Content Moderation?
@@
CCS were created with the stated goal of addressing the misinformation problem (IntroducingBirdwatchCommunitybaseda), but it is unclear if the platforms implementing them are serious about this aim. For example, many see Meta’s choice to replace their U.S. fact-checking program with a CCS (heatheraMoreSpeechFewer2025) as a thinly-veiled effort to curry political favor with a new presidential administration that is critical of fact-checking (nixMetaEndsFactchecking2025). In fact, from a platform’s perspective, CCS have other appealing properties: they cost less than professional fact-checkers, they shield platforms from accusations of biased policy enforcement, and they offload the difficult responsibility of distinguishing fact from fiction onto users. It may be expedient for platforms to build CCS in the current moment, but it is unclear if such systems will survive if the political climate shifts. Still, as long as CCS are in a position to exert such influence over the information ecosystem, it is important for the HCI research community to study them. Most importantly, we emphasize the need to better understand how CCS perform compared to other content moderation systems. CCS may perform well in certain cases, but if on the whole they perform worse than the systems they are replacing, for example on the question of coverage, they should be used thoughtfully.
@@
While it has been suggested that they can, CCS cannot effectively replace a fuller spectrum of content moderation tools, and were not designed to do so (elliottElonMusksMain). Content moderation is a broad term that encompasses a variety of different goals, including preventing crimes, mitigating lawful but awful behavior, and promoting high quality content (grimmelmann2015virtues). The tools used as a part of content moderation are similarly varied and include both “hard” approaches, like content take-downs and account bans, and “soft” approaches, such as content labeling (zannettouWonElectionEmpirical2021). Systems that provide context are narrow in their their goal and mechanism: they are a soft moderation attempt to aid content interpretation by providing users with additional information (matamoros25). And yet, the empirical results summarized in this paper demonstrate that CCS may play a meaningful role in the content moderation ecosystem, and it is worth exploring how they may integrate with other parts of it. For example: how might CCS be integrated into platforms’ automated moderation systems? In the Platform Treatment design aspect, none of the implementations that we examined used the CCS output to affect the visibility of posts. Perhaps these systems can be integrated into platform content recommendations algorithms to promote or demote posts based on their notes.
8.2. Underexplored CCS Design Aspects
@@
Our design space reveals several ways forward for a HCI research agenda on CCS. Several aspects are underexplored by current research, particularly the Participation aspect. Encouraging positive participation in collaborative computing systems is a classic topic in HCI research (krautBuildingSuccessfulOnline2012) and one in which the community could make strong contributions. One dimension of participation which has not yet been explored is how platforms may play a more active role in soliciting notes and ratings from contributors. Past HCI work on how to best make use of the various skills of crowds may be instructive here (bernsteinSoylentWordProcessor2015; kitturFutureCrowdWork2013). It is worth noting that the more consequential these systems become, the more they will be targets for negative forms of participation as well, primarily gaming (danielQualityControlCrowdsourcing2018; epsteinWillCrowdGame2020). Bridging algorithms are designed to be resistant to certain types of gaming (wojcikBirdwatchCrowdWisdom2022b), but vulnerabilities may emerge. One other area for study is the Presentation aspect, specifically how different approaches to visually displaying notes may affect user understanding of misleading claims (gamageLabelingSyntheticContent2025; martelMisinformationWarningLabels2023). Another promising area is the Inputs aspect, where recent work has begun to explore the ways in which LLMs can supplement human effort and increase the quality and quantity of notes (supernotes; mohammadiAIFeedbackEnhances2025a). Of particular interest is X’s recent decision to allow AI note contributors via a new API (liScalingHumanJudgment2025a). How such a decision affects the quantity and quality of notes on the platform should be a near-term focus of HCI research. Of course, studying how changes to CCS impact information dynamics across a platform will be difficult to study without more innovations in the Transparency aspect. X’s approach to transparency in their Community Notes system has been a boon for studying CCS. Hopefully other platforms will soon adopt similar approaches to support ongoing studies of CCS effectiveness.
8.3. Normative Guidelines for CCS Design
CCS are complex sociotechnical systems (chernsPrinciplesSociotechnicalDesign1976) that integrate people, algorithms, and speech. These systems’ growing influence over the information shown on major social media platforms means that the design choices made in their implementation can have significant social consequences (friedmanBiasComputerSystems1996; flanaganEmbodyingValuesTechnology2008). With such high stakes, it is crucial that a human-centered research agenda consider the implications of different design choices, in order to holistically evaluate these systems’ impacts. One way to do so is to identify the primary normative concerns of such a system. We emphasize three interconnected concerns that should guide CCS design: user informedness, distribution of power, and fairness.
8.3.1. User Informedness
While social media serves many different purposes, CCS are specifically concerned with social media’s ability to inform the public. Such systems should thus be evaluated by how well they facilitate user informedness: for example, the degree to which they inform users about the accuracy of information, minimize the spread of misleading posts, or otherwise improve information quality on a platform. Our literature review surfaced ample evidence that the notes produced by CCS do have the power to increase user informedness, both by decreasing an individual’s belief in misleading information, and by decreasing the spread of this information (wojcikBirdwatchCrowdWisdom2022b; renaultCollaborativelyAddingContext2024a; slaughter25; chuaiCommunitybasedFactcheckingReduces2024a; rollout21; diffusion21). Additionally, there is evidence (kimDifferentialImpactIndividual2025e) that these notes may produce less of a backfiring effect than forms of social correction that is not mediated by a CCS (mosleh21). But while published notes have been shown to be effective at promoting understanding, we must also question how effective CCS are at producing these notes. Several of these same studies found that Twitter’s implementation of CCS produces notes too slowly to prevent the spread of the most viral content (rollout21; chuaiCommunitybasedFactcheckingReduces2024a; renaultCollaborativelyAddingContext2024a). There is also a significant gap in the research on the question of coverage, or how many of the misleading posts on a platform receive “helpful” notes, which is currently unanswered. The research suggests that there is a CCS design trade-off between speed/coverage and quality/helpfulness: X’s implementation has optimized for the latter, perhaps at the expense of the former. Whether this is the right decision should be up for debate.
For designers of CCS hoping to improve user informedness by increasing the system’s speed and coverage, there are several aspects from our design space that could be relevant. For example, within the curation aspect, designers could explore bridging algorithm changes that could lead to faster identification of “helpful” notes. Additionally, design changes within the participation and inputs aspects, such as those that would allow a larger contributor base or that would actively solicit notes on the most viral posts, could lead to a higher volume of notes and ratings, and thus quicker selection of helpful notes. When it comes to the difficulty of assessing the coverage of CCS, design changes within the transparency aspect could make this easier to study, perhaps by providing easier access to post engagement metrics.
8.3.2. Distribution of Power
One of the selling points of CCS is that they shift power over content moderation from the platform to the community. Giving users more of a say over the content that they see on their platform can make moderation decisions more legitimate and trusted (shusas24; liu23; drolsbachCommunityNotesIncrease2024a), which is an admirable goal. Yet such claims require further scrutiny, and CCS should be evaluated to the extent that they actually empower users and do not merely allow platforms to shirk responsibility for misleading content. Evidence from our literature review suggests that bridging-based algorithms, by design, fail to produce notes for the most polarizing content (bouchaudAlgorithmicResolutionCrowdsourced2025b). For example, X’s implementation of the bridging algorithm requires polarized contributors to agree that a note is helpful in order to publish it. The editorial function of the bridging algorithm can thus limit the type and scope of contributions in two important ways: in terms of framing, the algorithm may prefer centrist points of view and content that could be agreeable to both sides of an issue; in terms of selection, agreement on a note may not occur for contentious topics, preventing its publication. Whether or not users at the ends of the ideological spectrum should be empowered to make moderation decisions is a valid question, but it is uncontroversial that the most polarizing content needs moderation, too. Research has found evidence of harmful content being handled poorly by the system, either because potential notes never reach helpful status (chujyo25), or because notes that do are not actually helpful for certain types of content, such as disinformation that is couched in humor (matamoros25). Zooming out, the low publishing rate for suggested notes (around  10% on X (Wirtschafter_Majumder_2023) and  6% on Meta [^8] https://x.com/guyro/status/1965800300557857053) gives us pause that the system is actually empowering these note authors. A related, underexplored area for research involves studying which note contributors are empowered by CCS. This research area is difficult to study holistically, as these systems are pseudonymous, but some work has begun to explore it qualitatively (yoon25) and via partnerships with platforms (allen22). More research is needed to understand how CCS can empower users of certain demographics or political views over others.
Our design space offers several ways that designers can further empower users via CCS. First and foremost is innovation in the participation aspect that would increase the set and diversity of contributors. The challenge for the platforms will be ensuing that quality remains as the participating contributor base increases, which is a tension that has been thoroughly studied in the crowdsourcing literature (danielQualityControlCrowdsourcing2018), for example. Innovation in the inputs aspect may increase the quality of notes across the board, perhaps by giving note authors helpful feedback (mohammadiAIFeedbackEnhances2025a), empowering more users to make successful contributions to the system. Another route to shift power to users is to disempower the platform. Though CCS are often described as ways to reduce platforms’ centrality in content moderation, platforms still exert power over the system in important ways, such as by deciding who is admitted as a contributor. Platforms could further guarantee their neutrality through changes to the transparency aspect, such as by making their contributor selection process public. But considering that X’s CCS implementation is the only even partially open system that we examined, a first step is for other platforms to follow X’s lead and open source their note curation algorithm and the crowdsourced data that they collect.
8.3.3. Fairness
CCS are often touted for their potential to implement content moderation in an unbiased way. Yet these systems are not inherently free from bias. Results from our literature review detail bias in contributor behavior, in terms of the types of accounts that they leave notes on, as well as how they rate suggested notes (allen22). Bridging algorithms have helped mitigate bias in the note curation process, with the caveats noted above, but additional bias can still be introduced—for example, in the note creation process, or in the selection of contributors. To give a concrete example: X’s bridging algorithm finds agreement among contributors, not the entire user base. If contributors’ beliefs are not representative of users’ beliefs then the algorithm’s outcomes may be a biased representation of the community, and not necessarily have broad appeal among users. There is some evidence that so far X’s implementation has not suffered from this problem, but it is worth considering. Also, accusations of bias in top-down moderation are often due to the opacity of such approaches and may be mitigated by transparency efforts. Only X’s implementation has provided any transparency, and yet the lack of transparency about the contributor base has made analyzing them for bias impossible. Finally, the approach of displaying context rather than removing posts is arguably fairer in that it lessens the damage from false positives, by empowering users to come to their own conclusions about content’s veracity (pengRescuingCounterspeechBridgingBased2024a). However, there is a trade off: leaving up questionable content, even with helpful context, can potentially do harm to those targeted by the content.
Designers should consider several aspects from our design space that can influence the fairness properties of a CCS. First off, all platforms have taken the approach that it is fairer not to remove misleading content. However designers may explore other options from the platform treatment dimension to identify certain cases where the platform should take action on content flagged by the crowd as misleading. Second, current implementations tend to display a single, definitive note, even if there may be several that are determined to be “helpful”. The presentation aspect offers alternatives that could further the aim of allowing users to come to their own conclusions, such as by showing multiple notes when there are multiple with helpful ratings. Additionally showing information about the note author could be helpful, as it could help users identify the credibility of the source. This may be at odds with contributor privacy, which current implementations have opted to protect. Finally, innovations in the curation aspect could explore other methods of selecting notes to display that may be regarded as fairer. Current implementations of the bridging algorithm use past ratings to identify contributors with diverse perspectives: perhaps it would be fairer to instead use in-app behavior? This could provide a fuller picture of a contributor’s ideological leanings. Encouraging more transparency would also give assurances of fairness, both in how contributors are selected, and how notes are published.
8.4. Limitations
Since there was no agreed upon name for CCS before our paper, we decided to focus our literature review on studies of Twitter-style CCS, and thus may have missed earlier work that would have meaningfully informed the design space. Our analysis of platforms’ implementations of CCS is based on public documentation and may not perfectly reflect the actual features of the deployed systems. Finally, our design space includes several non-exhaustive, speculative options that were not seen in any implementations, but nonetheless are included to illuminate alternative possibilities.
@@ 9. Conclusion
Crowdsourced Context Systems are increasingly prevalent in the content moderation strategies of prominent social media platforms. To understand these systems, their impacts, and how they may be better designed, we conducted a systematic literature review of existing research studying implementations of CCS and analyzed several major social media platforms’ implementations of such systems. Despite recent trends towards the adoption of CCS, we emphasize that they are only one piece of a larger content moderation suite, and suggest against using them as drop-in replacements for traditional fact-checking approaches. The framework we produced can be used to reason about these systems, and consists of a theoretical model and a design space. We additionally discuss some of the key normative concerns that should be considered when designing CCS. Such systems have potential to address the challenge of misinformation, but to realize this potential, future design research from the HCI community must focus on ways to understand and improve these systems’ performance.